a) i. The U.S. uses roughly 4.0 million Gigawatt-hours per year in electricity. Currently 50% of our electricity comes from coal. If every 1.0 kilowatt-hour of coal burned releases 2 lbs. of carbon dioxide, how much carbon dioxide is released from burning coal for electricity each year in the US? ii. What will be the reduction of carbon dioxide entering the atmosphere each year if coal usage could be reduced by 15% by utilizing solar energy? iii. List and discuss TWO advantages to using wind energy to generate electricity.
Answer #1
  • first we should know that for every 1 pound of coal , same amount of electricity is produced. viz
  • 1 pound of coal = 1 Kwh of electricity produced
  • so, if we say every 1.0 kilowatt-hour of coal burned releases 2 lbs. of carbon dioxide, it also means 1 pound of coal is being burned.

so, 4.0 million Gwh of electricity is used up every year in the US. which is = 4*106* 106 Kwh

= 4 *1012 Kwh

its 50% of electricity production comes from coal power. so that will be = 50% of 4 *1012 Kwh

= 2*1012 Kwh

so , if 1 Kwh of coal ( or 1 pound of coal) burned releases 2 pounds of Carbon dioxide in the atmosphere,

then amount of CO2 released by burning  2*1012 Kwh of coal = 2 x 2*1012

= 4 x 1012 pounds of CO2

or  = 4 trillion pounds.

B) the reduction of carbon dioxide entering the atmosphere each year if coal usage could be reduced by 15%

= 0.85 x(2*1012 Kwh) x 2 lbs

= 3.4 *1012 lbs of CO2

reduction in the amount of carbon dioxide => ( 4 x 1012 - 3.4 x 1012) = 6x1011 pounds of CO2

or = 0.6 trillion pounds of CO2

C) the advantages of using wind energy to produce electricity are:

i) it is a renewable and sustainable source of energy which is available ample is quantity, and moreover it is eco friendly source of electricity production , since it does not emit any greenhouse gases. as wind turbines are dependent only upon the wind to produce power.

ii) though wind turbine power generation capacity is not as high as coal and petroleum powered electricity generation but even though it can generate 3-5 million kWh of electricity in a year and its also cheap and easy to maintenance.